# Sequential Attention-based Sampling for Histopathological Analysis

### v1

@@ 3.4 Sequential Attention-based Sampling for Histopathological Analysis
During each training episode, at each timestep $t$, the agent observes the WSI state $S_{t}$, which is initialized with the low-resolution feature $S_{0}=Z$. Based on $S_{t}$, the policy network $\pi_{\theta}$ produces a probability distribution over a set of actions from a discrete action space $\{1,2,\ldots,N\}$. The agent samples an action $a_{t}$, corresponding to the next low-resolution patch to zoom in, according to the current policy $\pi_{\theta_{t}}(\cdot\mid S_{t})$ with a stochastic sampling rule. The zoomed-in patch is passed through HAFED to obtain its high-resolution representation $V(a_{t})$. This representation is then used by the TSU to selectively refine the current state $S_{t}$, by modifying only those patches that are similar to $a_{t}$ in the feature space. This produces an updated state $S_{t+1}$. The classifier trained together with HAFED then uses the updated state $S_{t+1}$ to produce a slide-level prediction $\hat{y}_{t}$. This prediction is compared against the ground truth label $y$ to compute a time-step wise reward, based on a negative binary cross-entropy (BCE) loss. This feedback is used to update the policy network using Proximal Policy Optimization (PPO) [28] algorithm. We employed Generalized Advantage Estimation (GAE) in computing advantage function [27], whose output was trained to predict the reward until the end of the episode. Both the policy and value network were modeled on standard architectures [14] details are in Appendix B.2.3. The same sequence of steps was followed during evaluation, except that the action with the highest probability based on the policy network’s output was chosen deterministically.
@@ 4.1 Dataset and implementation details
We evaluate the effectiveness of SASHA with two publicly available benchmark datasets CAMELYON16 breast cancer [3] and TCGA-NSCLC lung cancer dataset [23] dataset. While the CAMELYON16 labels correspond to normal versus pathological WSIs, TCGA-NSCLC labels correspond to distinguishing two different types of cancers, adenocarcinoma versus squamous cell carcinoma. We followed standardized train-test splits for both datasets. Further details on the datasets, hyperparameter selection, implementation, and training are provided in Appendix B.1 and B.2.
@@ 4.2 Baselines and Metrics
For a fairer comparison, we evaluate the performance of models that sample only a fraction of the WSI. Specifically, we evaluate RLogist trained to sample either exactly $10\%$ or exactly $20\%$ of patches at high resolution and compare it with our model trained under identical settings; this analysis also enables analyzing the impact of observation budgets on model performance. In addition, we also compare two other sota attention-based models ACMIL and DTFD, but evaluate their performance under same set of observations budgets, with random sampling of high-resolution patches (Table 2); details in Appendix B.3. In all cases, models were trained with the official code provided by the respective studies, except for RLogist. The official RLogist codebase lacks major elements, including state update models, possibly due to proprietarity; in this case, we report the performance metrics as reported in the original study on the same benchmarks [36].
@@ 4.3 Classification Performance, Inference Time and Compressibility
WSI compressibility. Similarly, a comparison of the WSI compressibility revealed a significant (>16×>16$\times$> 16 ×) improvement, compared to models that examine the entire slide at high resolution (Fig. 3d). The main reason for this saving is the fact that our HAFED model encodes features for both low- and high-resolution patches in a uniform $N\times d$ dimensional space; by contrast, other models encode high-resolution features in a $N\times k\times d$ dimensional space, where $k$ reflects the ratio of the zoomed-in to the scanning resolution size. The full set of results, including a comparison of model sizes, is shown in Appendix B.4. These results highlight the efficiency of our method in balancing performance with computational and representational costs.
@@ 4.4 Ablation Study and Limited Sampling
We performed a systematic ablation study removing each component of the model in turn – or replacing it with a naive variant – and evaluating it with the CAMELYON16 dataset. These included: a) replacing the WSI-pretrained ViT with a ResNet50, used commonly for feature extraction [37, 36], b) replacing multi-branch attention with single branch in the Classifer [14], c) replacing the targeted state update policy with a global state update [36], d) removing the TSU and updating the state only of the sampled (local, high-resolution) patch, and e) and selecting an action based on a random policy (see “Variant”s in Table 2). In every case, we observed sharp drops in accuracy, ranging from 7-45%, relative to the baseline SASHA model; similar drops occurred in AUROC and F1 scores Table 2. The strongest impact occurred upon changing the RL policy (Table 2, e), with the next most impactful effects occurring upon changing the feature extractor (a) or the state update method (c-d). The least impact occurred upon changing the classifier to single branch attention (b), but even here, the F1 score dropped by 14% points. Additional ablation experiments, and including exploring the effect of terminal versus intermediate reward during RL agent training, adopting a stochastic action selection policy are described in Appendix B.3. These results indicate that every component of our model was critical for achieving high performance, comparable to sota models.
@@ 5 Conclusions and Limitations
We show that an attention-augmented deep RL agent outperforms or performs on par with sota models on WSI classification, albeit at a fraction of their memory and computational costs. Key limitations of the model include a significant impact of the observation budget on model calibration, and the need to train the model with all high resolution patches, which significantly increases training time; other limitations are discussed in Appendix B.6. Nevertheless, we demonstrate explainable, reliable, and calibrated predictions, which could enhance clinicians’ trust in model predictions.
@@ A.2.1 Patch feature distillation with HAFED
During inference, at each time step, the first stage of HAFED (FA), receives as input $k$ zoomed-in sub-patch images corresponding to the selected action. It assigns attention scores to estimate the relative importance of each sub-patch and aggregates their features by computing a weighted sum based on these scores; this is then used to update the state (see next subsection, Section B.2.2). In the final time step, the second stage of HAFED (classifier), takes the state representation $S_{T}$ and predicts the label $\hat{y}$. In addition to the final classification, the second stage is also used to assign rewards during SASHA’ RL agent training (see subsection Section B.2.3).
@@ A.2.3 Actor and Critic
During inference, given the current state $S_{t}$, the Actor network outputs a policy in the form of a categorical probability distribution over the $N$ patches. The action $a_{t}$ is then selected deterministically as the patch index with the highest probability. In a later section (Section B.3.2) we explore the effect of alternative, stochastic selection policies during inference.
@@ A.3.2 Alternative Action Policies during Inference
In the current SASHA implementation, the inference strategy follows a deterministic, greedy policy wherein, for a given state, the action with the highest probability is always selected. This means the agent always picks the most likely action as per the learned policy distribution during inference (but not during training, see Section B.2.3).
@@ A.3.3 Alternative Reward Schedules during Training
The default SASHA model employs a step-wise reward during training, where the agent receives a reward at every time step of the episode associated with the label prediction loss at that time step (see Section B.2.3). In contrast, a previous competing model (RLogist [36]) employed a terminal reward strategy, providing feedback only at the end of each episode.
@@ A.6 Limitations
Appendix B Appendix
B.1 Dataset details
Dataset Task Train Val Test
Class 1 Class 2 Class 1 Class 2 Class 1 Class 2
CAMELYON16 Classify tumor 140 102 18 9 80 49
TCGA-NSCLC Classify sub-type 339 319 86 81 104 109
B.2 Implementation and Hyper-parameter Selection
B.2.1 Patch feature distillation with HAFED
Layer name Input Dim Output Dim Input Batch Output Batch
$V_{i}={\alpha_{1}}^{i}\cdot U_{i},\quad V_{i}\in\mathbb{R}^{d}$
Layer-4 384 128 $V=\{V_{1},V_{2},\ldots,V_{N}\}$ $C=\{c_{1},c_{2},\ldots,c_{N}\}$
Layer-5 384 128 $V=\{V_{1},V_{2},\ldots,V_{N}\}$ $D=\{d_{1},d_{2},\ldots,d_{N}\}$
Layer-7 384 2 $h$ logits
During inference, at each time step, the first stage of HAFED (FA), receives as input $k$ zoomed-in sub-patch images corresponding to the selected action. It assigns attention scores to estimate the relative importance of each sub-patch and aggregates their features by computing a weighted sum based on these scores; this is then used to update the state (see next subsection, Section B.2.2). In the final time step, the second stage of HAFED (classifier), takes the state representation $S_{T}$ and predicts the label $\hat{y}$. In addition to the final classification, the second stage is also used to assign rewards during SASHA’ RL agent training (see subsection Section B.2.3).
B.2.2 Updating WSI state with TSU
Layer Input Dim Output Dim # Parameters
FC-1 1152 768 885,504
FC-2 768 384 295,296
Layer Norm - 384 384
B.2.3 Actor and Critic
[…]
During inference, given the current state $S_{t}$, the Actor network outputs a policy in the form of a categorical probability distribution over the $N$ patches. The action $a_{t}$ is then selected deterministically as the patch index with the highest probability. In a later section (Section B.3.2) we explore the effect of alternative, stochastic selection policies during inference.
Model Hyper-parameters Selection Criterion SASHA - 0.1 / 0.2
CAMELYON16 TCGA-NSCLC
Range Values Range Values
HAFED Attn-2 in HAFED (M) Validation 1-5 5 1-5 1
Attn -1 Masking Probability Validation 0.4 - 0.6 0.4 0.4 - 0.6 0.4
Attn -2 Masking Probability Validation 0.4 - 0.6 0.4 0.4 - 0.6 0.4
Actor & Critic Discount Factor Validation 0.8-1.0 1.0 0.8-1.0 1.0
Clipping Parameter Default - 0.1 - 0.1
Entropy Regularization Coefficient Default - 0.001 - 0.001
Gradient Clipping Threshold Default - 0.5 - 0.5
B.3 Additional Control Experiments
B.3.1 Alternative Models under an Observation Budget
Model Accuracy AUC F1
ACMIL - 0.1 0.841$\pm$0.036 0.828$\pm$0.034 0.773$\pm$0.038
DTFD - 0.1 0.829$\pm$0.013 0.835$\pm$0.039 0.829$\pm$0.013
SASHA - 0.1 0.901$\pm$0.021 0.918$\pm$0.014 0.856$\pm$0.031
ACMIL - 0.2 0.838$\pm$0.035 0.857$\pm$0.030 0.773$\pm$0.043
DTFD - 0.2 0.862$\pm$0.027 0.878$\pm$0.030 0.862$\pm$0.027
SASHA - 0.2 0.953$\pm$0.017 0.979$\pm$0.008 0.937$\pm$0.024
B.3.2 Alternative Action Policies during Inference
In the current SASHA implementation, the inference strategy follows a deterministic, greedy policy wherein, for a given state, the action with the highest probability is always selected. This means the agent always picks the most likely action as per the learned policy distribution during inference (but not during training, see Section B.2.3).
Inference Policies Accuracy AUC F1
Stochastic - full distribution 0.884 0.846 0.831
Stochastic - top-$k$ ($k=3)$ 0.876 0.826 0.814
Stochastic - top-$k$ ($k=5)$ 0.868 0.823 0.805
Stochastic - top-$p$ ($p=0.9)$ 0.876 0.859 0.818
Stochastic - top-$p$ ($p=0.8)$ 0.868 0.851 0.795
Deterministic (SASHA default) 0.953 0.979 0.937
B.3.3 Alternative Reward Schedules during Training
The default SASHA model employs a step-wise reward during training, where the agent receives a reward at every time step of the episode associated with the label prediction loss at that time step (see Section B.2.3). In contrast, a previous competing model (RLogist [36]) employed a terminal reward strategy, providing feedback only at the end of each episode.
Last time step (Terminal) 0.892 0.941 0.862 15.5 0.0020
Each time step (SASHA default) 0.953 0.979 0.937 24.25 0.0032
B.3.4 Analysis of RL Agent Behavior
B.4 Computation of WSI compressability, Inference time and ECE
The compressibility factor $C$ is defined as:
  • Accuracy: the proportion of correctly classified samples in the bin.
  • Confidence: the average of the predicted probabilities in the bin.
$\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i})$
$\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$
B.5 Hardware
B.6 Limitations
[…]

### v2

@@ 3.4 Sequential Attention-based Sampling for Histopathological Analysis
During each training episode, at each timestep $t$, the agent observes the WSI state $S_{t}$, which is initialized with the low-resolution feature $S_{0}=Z$. Based on $S_{t}$, the policy network $\pi_{\theta}$ produces a probability distribution over a set of actions from a discrete action space $\{1,2,\ldots,N\}$. The agent samples an action $a_{t}$, corresponding to the next low-resolution patch to zoom in, according to the current policy $\pi_{\theta_{t}}(\cdot\mid S_{t})$ with a stochastic sampling rule. The zoomed-in patch is passed through HAFED to obtain its high-resolution representation $V(a_{t})$. This representation is then used by the TSU to selectively refine the current state $S_{t}$, by modifying only those patches that are similar to $a_{t}$ in the feature space. This produces an updated state $S_{t+1}$. The classifier trained together with HAFED then uses the updated state $S_{t+1}$ to produce a slide-level prediction $\hat{y}_{t}$. This prediction is compared against the ground truth label $y$ to compute a time-step wise reward, based on a negative binary cross-entropy (BCE) loss. This feedback is used to update the policy network using Proximal Policy Optimization (PPO) [28] algorithm. We employed Generalized Advantage Estimation (GAE) in computing advantage function [27], whose output was trained to predict the reward until the end of the episode. Both the policy and value network were modeled on standard architectures [14] details are in Appendix A.2.3. The same sequence of steps was followed during evaluation, except that the action with the highest probability based on the policy network’s output was chosen deterministically.
@@ 4.1 Dataset and implementation details
We evaluate the effectiveness of SASHA with two publicly available benchmark datasets CAMELYON16 breast cancer [3] and TCGA-NSCLC lung cancer dataset [23] dataset. While the CAMELYON16 labels correspond to normal versus pathological WSIs, TCGA-NSCLC labels correspond to distinguishing two different types of cancers, adenocarcinoma versus squamous cell carcinoma. We followed standardized train-test splits for both datasets. Further details on the datasets, hyperparameter selection, implementation, and training are provided in Appendix A.1 and A.2.
@@ 4.2 Baselines and Metrics
For a fairer comparison, we evaluate the performance of models that sample only a fraction of the WSI. Specifically, we evaluate RLogist trained to sample either exactly $10\%$ or exactly $20\%$ of patches at high resolution and compare it with our model trained under identical settings; this analysis also enables analyzing the impact of observation budgets on model performance. In addition, we also compare two other sota attention-based models ACMIL and DTFD, but evaluate their performance under same set of observations budgets, with random sampling of high-resolution patches (Table 2); details in Appendix A.3. In all cases, models were trained with the official code provided by the respective studies, except for RLogist. The official RLogist codebase lacks major elements, including state update models, possibly due to proprietarity; in this case, we report the performance metrics as reported in the original study on the same benchmarks [36].
@@ 4.3 Classification Performance, Inference Time and Compressibility
WSI compressibility. Similarly, a comparison of the WSI compressibility revealed a significant (>16×>16$\times$> 16 ×) improvement, compared to models that examine the entire slide at high resolution (Fig. 3d). The main reason for this saving is the fact that our HAFED model encodes features for both low- and high-resolution patches in a uniform $N\times d$ dimensional space; by contrast, other models encode high-resolution features in a $N\times k\times d$ dimensional space, where $k$ reflects the ratio of the zoomed-in to the scanning resolution size. The full set of results, including a comparison of model sizes, is shown in Appendix A.4. These results highlight the efficiency of our method in balancing performance with computational and representational costs.
@@ 4.4 Ablation Study and Limited Sampling
We performed a systematic ablation study removing each component of the model in turn – or replacing it with a naive variant – and evaluating it with the CAMELYON16 dataset. These included: a) replacing the WSI-pretrained ViT with a ResNet50, used commonly for feature extraction [37, 36], b) replacing multi-branch attention with single branch in the Classifer [14], c) replacing the targeted state update policy with a global state update [36], d) removing the TSU and updating the state only of the sampled (local, high-resolution) patch, and e) and selecting an action based on a random policy (see “Variant”s in Table 2). In every case, we observed sharp drops in accuracy, ranging from 7-45%, relative to the baseline SASHA model; similar drops occurred in AUROC and F1 scores Table 2. The strongest impact occurred upon changing the RL policy (Table 2, e), with the next most impactful effects occurring upon changing the feature extractor (a) or the state update method (c-d). The least impact occurred upon changing the classifier to single branch attention (b), but even here, the F1 score dropped by 14% points. Additional ablation experiments, and including exploring the effect of terminal versus intermediate reward during RL agent training, adopting a stochastic action selection policy are described in Appendix A.3. These results indicate that every component of our model was critical for achieving high performance, comparable to sota models.
@@ 5 Conclusions and Limitations
We show that an attention-augmented deep RL agent outperforms or performs on par with sota models on WSI classification, albeit at a fraction of their memory and computational costs. Key limitations of the model include a significant impact of the observation budget on model calibration, and the need to train the model with all high resolution patches, which significantly increases training time; other limitations are discussed in Appendix A.6. Nevertheless, we demonstrate explainable, reliable, and calibrated predictions, which could enhance clinicians’ trust in model predictions.
@@ A.2.1 Patch feature distillation with HAFED
During inference, at each time step, the first stage of HAFED (FA), receives as input $k$ zoomed-in sub-patch images corresponding to the selected action. It assigns attention scores to estimate the relative importance of each sub-patch and aggregates their features by computing a weighted sum based on these scores; this is then used to update the state (see next subsection, Section A.2.2). In the final time step, the second stage of HAFED (classifier), takes the state representation $S_{T}$ and predicts the label $\hat{y}$. In addition to the final classification, the second stage is also used to assign rewards during SASHA’ RL agent training (see subsection Section A.2.3).
@@ A.2.3 Actor and Critic
During inference, given the current state $S_{t}$, the Actor network outputs a policy in the form of a categorical probability distribution over the $N$ patches. The action $a_{t}$ is then selected deterministically as the patch index with the highest probability. In a later section (Section A.3.2) we explore the effect of alternative, stochastic selection policies during inference.
@@ A.3.2 Alternative Action Policies during Inference
In the current SASHA implementation, the inference strategy follows a deterministic, greedy policy wherein, for a given state, the action with the highest probability is always selected. This means the agent always picks the most likely action as per the learned policy distribution during inference (but not during training, see Section A.2.3).
@@ A.3.3 Alternative Reward Schedules during Training
The default SASHA model employs a step-wise reward during training, where the agent receives a reward at every time step of the episode associated with the label prediction loss at that time step (see Section A.2.3). In contrast, a previous competing model (RLogist [36]) employed a terminal reward strategy, providing feedback only at the end of each episode.